(c) Show that the value of V1 + cos x dx is less than 1.6.
Answer #1

{ Ces(24) = 2651-12 (Cesn=265716-1 Suitcasa de -2 J1 +265 4/6-1 de = S 12 Gos?lla dan = so Tes Ula 1 dhon 258 [1 sinala] - 25

answered by: ANURANJAN SARSAM
Answer #2

{ Ces(24) = 2651-12 (Cesn=265716-1 Suitcasa de -2 J1 +265 4/6-1 de = S 12 Gos?lla dan = so Tes Ula 1 dhon 258 [1 sinala] - 25